Jebediah:

A student is trying to solve the set of two equations given below: Equation A: x + z = 6 Equation B: 2x + 4z = 1 Which of the following is a possible step used in eliminating the z-term? Also, as I said before, could you explain your answer? Thanks guys >_<     Multiply equation B by 4. Multiply equation A by 2. Multiply equation A by –4. Multiply equation B by 2.

Shadow:

The way that elimination works is that you add the two equations together. However, before you do so, you want to make sure that by doing so, a variable is eliminated. Here we want to eliminate z. Due to like terms, only z terms will be able to add by each other. We have z on the top equation, and 4z on the bottom equation. In order to make 4z zero, we must subtract by 4z. So if we add -4z to positive 4z, we will get zero. In order to get -4z on the top equation, what must we do to the top equation.

Jebediah:

umm not quite sure :/

Shadow:

Example. 2a + b = 2 6a + 8b = 10 I want to eliminate 8b. So I multiply the top equation by -8 \[-8(2a + b = 2) \] \[-16a -8b = -16\] Now if I add the two equations, -8b and +8b cancel each other out.

Jebediah:

Yes so we are left with the b?

Jebediah:

from the 8s

Shadow:

No. \[8b + (-8b) = 0\]

Shadow:

Or \[8b - 8b = 0\] Same thing.

Shadow:

It's basically 8 of something minus 8 of the same thing.

Shadow:

Does that make sense to you?

Jebediah:

yes it does

Shadow:

So what can we do to Equation A: x + z = 6, in order to take out the 4z on the bottom equation.

Jebediah:

hmm

Jebediah:

well if we ultiply the 2 by 2

Jebediah:

by -2*

Jebediah:

we would get -4 right?

Jebediah:

and then the 4's cancel eachother out?

Jebediah:

so z = 1?

Jebediah:

@Shadow

Jebediah:

ar you there??

Mehek:

No that's incorrect Multiplying by -2 would cancel x out and Shadow said to cancel z out

Jebediah:

oh ok

Jebediah:

so what would I have to do?'

Mehek:

What could you multiply x + z = 6 with so that 4z in the second equation gets taken out?

Jebediah:

6?

Jebediah:

I am not good at math >_<

Mehek:

If you multiply the first equation by 6: x + z = 6 -> 6x + 6z = 36 And then your second equation is 2x + 4z = 1 If you add them together 6x + 6z = 36 +2x + 4z = 1


8x + 10z = 37 The z is still there it didn't get canceled out

Jebediah:

oh

Jebediah:

I'm not sure

Mehek:

It would be -4 -4x - 4z = -24 + 2x + 4z = 1


-2x + 0 = -23 Then you solve for x -2x = -23 -23/-2 = ?

Jebediah:

11.5?

Mehek:

Yes

Mehek:

Now in either equation you can plug 11.5 for x We'll use the first one x + z = 6 11.5 + z = 6 Solve for z

Jebediah:

-5.5?

Mehek:

Yep Your question doesn't require to solve for this but I'm just showing you for later use

Mehek:

So back to your question

Mehek:

What did we do to x + z = 6 to get rid of z

Jebediah:

multiplied by -4

Mehek:

So option C)

Jebediah:

ok thx :D